### v1

# Hardware-efficient Customized Keyword Spotting with Spectral-Temporal Graph Attentive Pooling and a hybrid Loss

Abstract
Most of the existing systems designed for keyword spotting (KWS) rely on a predefined set of keyword phrases. However, the ability to recognize customized keywords is crucial for tailoring interactions with intelligent devices. In this paper, we present a novel framework for customized KWS. This framework leverages the hardware-efficient LiCoNet architecture as the encoder, enhanced by a spectral-temporal pooling layer and a hybrid loss function to facilitate effective word embedding learning. The experimental results on a substantial internal dataset have demonstrated the distinct advantages of the proposed framework. LiCoNet performs at a similar level ($1.98\%$ FRR at $0.3$ FAs/Hr) to the computationally intensive Conformer, which requires 13x computational resources.
Index Terms—  Query-by-example Keyword Spotting, Conformer, LicoNet, Spectral-temporal Attentive Pooling, AAM, SoftTriplet
1 Introduction
A keyword spotting (KWS) system serves the purpose of detecting a predetermined keyword within a continuous real-time audio stream. This capability is pivotal in facilitating interactions between users and voice assistants. The introduction of a customized KWS system, which empowers users to define their own keywords, offers a substantial degree of flexibility and personalization in user experiences. However, this customization also presents significant challenges, such as the need for a small KWS memory footprint, minimizing latency, and handling user-defined keyword phrases that may not align with the training data distribution.
One approach to address these challenges involves the use of Query-by-Example (QbyE) techniques [1]. In this context, the KWS system utilizes audio samples of keywords provided by users to generate fixed-length embeddings. These embeddings are then employed to assess the similarity between test samples and the enrolled keywords within the embedding space, ultimately determining the presence of a keyword. The uppermost diagram in Figure 1 illustrates a broad framework for QbyE KWS. Within this framework, there is a pooling layer positioned after the encoder, which is responsible for creating an information-rich embedding. This embedding is subsequently supplied to a classifier to distinguish between sub-words or words.
In previous studies, transformers have found extensive use in encoder modeling because of their substantial modeling capabilities [2] [3]. Nevertheless, attention-based models are associated with significant computational demands and impose a high runtime memory burden when deployed on hardware. This characteristic makes them unsuitable for an always-on KWS system. The linearized convolution network (LiCoNet) as introduced in [4] for KWS modeling, offers excellent hardware efficiency while maintaining a high level of model effectiveness.
In this study, we introduce a LiCoNet-based, hardware-efficient framework for customized KWS modeling. We employ spectro-temporal graph attentive pooling (GAP) [5] to generate informative embeddings. This pooling layer demonstrates strong capability in comprehending the complex relationships within spectral-temporal data. During the training phase, we formulate a hybrid loss function that combines elements from the Additive Angular Margin (AAM) and SoftTriplet losses, which are widely employed in tasks such as face recognition [6] and speaker recognition [7]. The hybrid loss is crafted to enhance the distinctiveness of words and phonemes while simultaneously reducing the variability in learned embeddings attributed to speakers. Our experimental results, conducted on a substantial internal dataset, showcase the advantages of our proposed framework, which features GAP and the hybrid loss, for customized KWS. Notably, LiCoNet achieves performance levels similar to those of the computationally intensive Conformer, which requires $13$x computational resources.
2 Methodology
2.1 Encoder-decoder Architecture
The system architecture is illustrated in Figure 1. It adopts an encoder-decoder structure during the training phase. The encoder takes the acoustic feature of a word phrase as input and produces an embedding that is subsequently forwarded into the decoder for classification. The pooling layer serves as a dimensionality reduction technique to create a concise yet informative embedding. During the testing phase, a user enrolls in the system by providing a few samples of the customized keyword. Detection occurs by comparing the embedding of the testing speech within a sliding window against the enrolled samples.
Fig. 1: The customized KWS training framework.
2.2 Feature Encoder
2.2.1 ECAPA_TDNN
The entire training and testing process shares similarities with the speaker verification (SV) task. Consequently, we consider ECAPA_TDNN, a commonly used backbone model architecture for the SV task [8], as a potential choice for the acoustic feature encoder in this study. The ECAPA_TDNN model consists of a 1D convolution followed by three 1D SE-Res2Blocks, 1D convolution, attentive statistical pooling, and a fully connected (FC) layer. After each layer within the SE-Res2Block, we apply non-linear ReLU activation and batch normalization (BN). The embedding feature vectors are extracted from the FC layer.
2.2.2 Convolution-augmented Transformer (Conformer)
The Conformer architecture has proven its remarkable effectiveness within the sequence-to-sequence domain [9] and has achieved significant success in the realm of speech recognition tasks [10] [11] [12]. This architecture seamlessly integrates the capabilities of both convolutional and self-attention mechanisms, providing a flexible and exceptionally potent solution for learning feature representations from sequential data. Each Conformer block comprises four consecutive modules, including a feed-forward module, a self-attention module, a convolution module, and a second feed-forward module [9]. This Conformer-based encoder demonstrates the ability to leverage position-specific local features, facilitated by the convolution module, while simultaneously capturing content-based global interactions through the self-attention module.
2.2.3 Linearized Convolution Network (LicoNet)
LiCoNet represents a hardware-efficient architecture specifically designed for the KWS task, as detailed in [4]. This architecture is carefully crafted as a streaming convolution network, using equivalent linear operators to ensure efficient inference while preserving a high level of detection accuracy. Each LiCo-Block is structured as a bottleneck configuration composed of three 1D convolution layers. The initial layer employs streaming convolution with a kernel size greater than $1$, followed by two subsequent point-wise convolutions.
2.3 Feature Aggregator
Pooling plays an essential role in neural architectures, serving the purpose of distilling crucial insights from sequential data while preserving essential contextual details. In the context of our study, we investigate two distinct pooling strategies for word embedding learning.
Attentive Statistic Pooling (ASP) ASP combines the strengths of both statistical pooling and attention mechanisms [13]. Attention allows the model to dynamically weigh the importance of different elements along the temporal dimension, enabling the extraction of salient features that are crucial for the task at hand.
Spectral-temporal Graph Attentive Pooling (GAP) GAP has gained success in the field of speech and audio processing [5] [14]. It leverages the power of graph neural networks to comprehend complex relationships within spectral-temporal data. The spectral and temporal attention module comprises three graph attention blocks, each housing the graph attention network (GAT) and graph pooling. This configuration empowers the model to adapt the pooling procedure dynamically, facilitating the extraction of crucial features.
2.4 Loss Function
2.4.1 Additive Angular Margin (AAM)
The AAM loss is designed to enhance the discrimination power of neural networks by emphasizing the angular separation between class embeddings and is prevalent in the context of face recognition and feature embedding [6]. The loss function is defined as,
$\small\mathcal{L}_{aam}=-\mathop{log}\frac{e^{\textbf{\em s}\>cos(\theta_{y_{i}}+m)}}{e^{\textbf{\em s}\>cos(\theta_{y_{i}}+m)}+\sum_{j=1,j\neq y_{i}}^{C}e^{\textbf{\em s}\>cos(\theta_{j})}},$ (1)
where $\theta_{j}$ is the angle between the feature $\mathbf{x}_{i}\in\mathbb{R}^{d}$ and the weight $\mathbf{w}_{j}\in\mathbb{R}^{d}$. $\mathbf{w}_{j}$ denotes the $j$-th column of the weight $[\mathbf{w}_{1},\cdots,\mathbf{w}_{C}]\in\mathbb{R}^{d\times C}$ of the last fully-connected layer that maps $d$-dimensional embeddings to the logits. $C$ is the number of classes. s is a rescaling factor. An additive angular margin $m$ is applied for adjustment.
Fig. 2: DET curves of Conformer using various loss formulations (left) and of different encoders using the hybrid loss (right).
2.4.2 SoftTriplet
For customized KWS tasks, it’s important to note that the training and testing datasets have no overlap in data distribution. This uniqueness necessitates that the model possesses strong generalization capabilities. The QbyE system can hence be conceptualized as an optimization problem featuring triplet constraints. The primary objective is to minimize the distance between the embedding of an enrolled keyword and those of the same word phrase while simultaneously maximizing the distance between embeddings of different word phrases. To address this optimization goal and simultaneously account for intra-class variance—such as speaker variability or variations in speaking rates—we have chosen to adopt the SoftTriplet loss [15] as a word-level loss function,
$\small\mathcal{S}^{\prime}_{i,c}=\sum_{k}\frac{exp(\frac{1}{\gamma}\mathbf{x}_{i}^{\top}\mathbf{w}_{c}^{k})}{\sum_{k}exp(\frac{1}{\gamma}\mathbf{x}_{i}^{\top}\mathbf{w}_{c}^{k})}\mathbf{x}_{i}^{\top}\mathbf{w}_{c}^{k}$ (2)
$\small\mathcal{L}_{st}({\mathbf{x}}_{i})=-\mathop{log}\frac{exp(\lambda(\mathcal{S}^{\prime}_{i,y_{i}}-\delta))}{exp(\lambda(\mathcal{S}^{\prime}_{i,y_{i}}-\delta))+\sum_{j}exp(\lambda\mathcal{S}^{\prime}_{i,j})},$ (3)
where $\mathcal{S}^{\prime}_{i,c}$ is the similarity between feature $\mathbf{x}_{i}\in\mathbb{R}^{d}$ and the class $c$. $\mathbf{w}_{c}^{k}$ is the $k$-th center out of $K$ centers for the class $c$. $\delta$ is a predefined margin. $\lambda$ denotes a scaling factor.
2.4.3 Hybrid Loss
Phoneme Context Phonemes serve as the fundamental phonetic units that compose spoken words. Incorporating the context of phonemes into modeling offers a nuanced source of information for refining word embeddings. In our approach, as depicted in Fig. 1, we introduce a dedicated phoneme classifier into the training framework. The phoneme loss is computed by aggregating the frame-level AAM loss, applied to phoneme labels, across all frames.
Speaker Variability Acoustic variations related to individual speakers, such as differences in pitch, tone, or pronunciation, exert a significant influence on speech modeling. Existing approaches in QbyE KWS often assume that the system user is the same as the enrolled speaker. To address and disentangle speaker dependency within the application, we have incorporated a reverse speaker loss into our methodology, with the objective of learning speaker-independent embeddings (see Fig. 1). More specifically, we have devised an AAM-based reverse speaker loss, which is employed to maximize the speaker classification loss through the application of a gradient reversal layer (GRL) [16] during the training process. The parameter-free GRL functions as an identity transform during forward propagation but reverses gradients during back-propagation, feeding them into the preceding layer.
Consequently, our hybrid loss function is constructed as a combination of word-level loss, phoneme-level loss, and the reverse speaker loss.
$\small\begin{split}\mathcal{L}(\mathbf{x},\mathbf{y})=\mathcal{L}_{st}\left(\mathbf{x},y^{w}\right)-\eta\mathcal{L}_{aam}(\mathbf{x},y^{s})+\mu\mathcal{L}_{aam}(\mathbf{x},y^{p}),\end{split}\vspace{-2ex}$ (4)
where $\mathbf{x}$ is the acoustic feature vector, $\mathbf{y}=(y^{w},y^{s},y^{p})$ $y^{w}$ is the word label, $y^{s}$ is the speaker label, $y^{p}\in\mathbb{R}^{T}$ is the phoneme label sequence, $\eta$ and $\mu$ are scaling factors.
3 Experiments
3.1 Dataset
We use the Librispeech [17] dataset containing 960 hours of read English audiobooks sampled at $16$ kHz along with transcriptions. We employ a pre-trained acoustic model for the force-alignment to segment utterances into individual words. Each word-level segment is standardized to $2$s long by clipping or zero padding on both sides of the audio.
We use the internal aggregated and de-identified keyword dataset for evaluation. The positive data contains $275.7$k utterances from $629$ speakers. The total duration of negative data is up to $200$ hours. We extract acoustic features using $40$-dimensional log Mel-filterbank energies computed over a $25$ms window every $10$ms.
[…]
Encoder Single Loss Hybrid Loss (Word AAM) Hybrid Loss (Word SoftTriplet)
Word CE
Word AAM
Speaker
Speaker + Phoneme
Speaker + Phoneme
Speaker + Phoneme
ECAPA_TDNN $16.28$ $12.09$ $10.81$ $8.95$ $7.29$ $5.58$
Conformer $11.98$ $8.26$ $4.88$ $4.77$ $2.33$ $\mathbf{1.63}$
LiCoNet $13.20$ $9.75$ $7.49$ $5.36$ $3.63$ $\mathbf{1.98}$
3.2 Experimental Setup
Model architecture We conduct the experiments on three model types: ECAPA_TDNN, Conformer, and LicoNet. We setup ECAPA_TDNN with $128$ channels in the convolution layers and a $64$ dimensional bottleneck in the SE-Block and attention module. The scale dimension $s$ in the Res2Block is $8$. Conformer has two heads per multi-headed self-attention layer with $128$ input and output nodes [18]. The linear hidden units have a dimensionality of $192$, and the convolution module uses the kernel size of $7$. We construct LiCoNet by stacking $5$ LiCo-Blocks with the expansion factor of $6$ and the kernel size of $5$ [4]. Table 2 presents the model size and floating point operations per second (FLOPs) of $2$s audio for each encoder model.
Feature aggregator We compare Graph Attentive Pooling (GAP) against Attentive Statistic Pooling (ASP) as the feature aggregator. The spectral, temporal and spectro-temporal attention blocks use pooling ratios of $0.71$, $0.86$, and $0.71$, respectively.
Loss function We focus on investigating the effectiveness of Additive Angular Margin (AAM) and SoftTriplet in customized KWS modeling, with cross-entropy as the baseline. The AAM Softmax margin $m$ and scale s in Eq. 1 are set to $0.2$ and $32$, respectively. The SoftTriplet loss uses the scaling factor $\lambda=60$, $\gamma=1$, the margin $\delta=0.03$, and $K=10$. $\eta$ and $\mu$ in Eq. 4 is set to $0.1$ and $0.5$, respectively.
Training and testing protocols All KWS models are trained to predict 1002 targets (i.e., the top $1$k frequent words, Silence, and unknown words). We use a batch size of 64 with 8 GPUs for 40-epoch training. We adopt the triangular2 policy as described in [19] in conjunction with the Adam optimizer with a cyclical learning rate increased from $1$e-$8$ to $1$e-$3$ in $20$k warming-up updates. During testing, $3$ utterances were randomly picked as enrollments. For a given query, the cosine distance is used to compare the similarity between the query embedding and the 3 enrolled ones. The minimum distance is used to compare against a threshold value to make the detection decision. We present the model performance by plotting detection error trade-off (DET) curves, where the x-axis and y-axis represent the number of false accepts (FA) per hour and false reject rate (FRR), respectively.
4 Results and Discussion
Model Performance
Table 1 summarizes FRR of different models at $0.3$ FAs/Hr. In the single word loss configuration, the AAM loss significantly outperforms the CE loss across different encoders. Specifically, AAM improves FRR by $25.7\%$ for ECAPA_TDNN, $31\%$ for Conformer, and $26.1\%$ for LiCoNet. In the hybrid loss configuration featuring the word AAM loss, the inclusion of the reverse speaker loss greatly decreases FRR, particularly for Conformer, resulting in a reduction of $40.9\%$. By incorporating the phoneme loss, we can notice additional enhancements. The efficacy of the hybrid loss underscores the value of using complementary information from both auxiliary losses for KWS modeling.
Table 2: Model size and computation cost of each encoder.
Encoder ##\##Params FLOPs
ECAPA_TDNN $540.1$K $39.1$M
Conformer $1.4$M $642.2$M
LicoNet $694.1$K $46.5$M
As for the feature aggregator, GAP delivers further substantial improvements compared to ASP across all encoders. In particular, FRR has been decreased by $18.5\%$ for ECAPA_TDNN, $51.1\%$ for Conformer, and $32.2\%$ for LiCoNet. These improvements align with the enhancements observed in the speaker verification task [5] and can be attributed to the increased discriminative capability introduced by the graph pooling strategy. Furhtermore, the SoftTriplet loss effectively captures potential unseen intra-variance within the evaluation data. In the hybrid loss configuration employing graph pooling, the word SoftTriplet loss consistently leads to the best system performance across all models, with a particularly impressive $45.4\%$ reduction in FRR for LiCoNet.
It is noteworthy to see that Conformer consistently maintains superior performance across various loss formulations and pooling strategies. However, LiCoNet achieves comparable performance to Conformer when utilizing the hybrid loss incorporating the word SoftTriplet loss and graph pooling strategy. In Fig. 2, we present DET curves of Conformer using various loss formulations and pooling strategies (left), and those of different encoders while using the hybrid loss featuring SoftTriplet and GAP (right).
Model Efficiency
As shown in Table 2, Conformer boasts the largest model size with considerably more computational demands, despite its superiority in terms of high model capacity. Conversely, LiCoNet strikes a favorable balance between model efficiency and effectiveness, offering performance on par with Conformer while keeping computational costs close to that of ECAPA_TDNN.
5 Conclusion
In this study, we introduce a hardware-efficient customized KWS system that is centered around the LiCoNet architecture and complemented by a spectral-temporal graph pooling layer and a hybrid loss function. The experimental results showcase the advantages of our framework featuring GAP and the hybrid loss: LiCoNet achieves performance levels similar to those of the computationally intensive Conformer.